# Novel thermodynamic inequality for charged and rotating AdS black holes

### v1

@@ Abstract
We propose a new thermodynamic inequality for stationary and asymptotically Anti-de Sitter charged and rotating black holes, $4\pi J^{2}/(3MV)<1$. This inequality is derived through an analysis of the roots of the identity which holds among the thermodynamic variables to avoid a naked singularity. We have analyzed the Kerr-AdS black hole and the asymptotically AdS uncharged and rotating black strings and find strong supporting evidence for different horizon topologies. Using this inequality we verify the validity of reverse isoperimetric inequality in its conventional form (${\cal R}\geq 1$) in the presence of rotation. By examining a wide range of black hole solutions, we confirm the validity of the proposed inequality. Assuming that reverse isoperimetric inequality continues to hold in higher dimensions, and employing the intermediate refined reverse isoperimetric inequalities introduced in [Phys. Rev. Lett. 131, 241401] as guiding principles, we extend the inequality to higher-dimensional spacetimes.
@@
where $T$, $V$, $\Omega$, and $\Phi$ are the Hawking temperature, thermodynamic volume, angular velocity, and electric potential respectively, we clearly observe that the mass $M$ of the KN–AdS black hole should be essentially identified as enthalpy rather than internal energy. Then, it is straightforward to have

### v2

# Novel thermodynamic inequality for rotating AdS black holes

@@ Abstract
We propose a new thermodynamic inequality for stationary and asymptotically Anti-de Sitter rotating black holes, $4\pi J^{2}/(3MV)<1$. This inequality is derived through an analysis of the roots of the identity which holds among the thermodynamic variables to avoid a naked singularity. We have analyzed the Kerr-AdS black hole and the asymptotically AdS uncharged and rotating black strings and find strong supporting evidence for different horizon topologies. Using this inequality we verify the validity of reverse isoperimetric inequality in its conventional form (${\cal R}\geq 1$) in the presence of rotation. By examining a wide range of black hole solutions, we confirm the validity of the proposed inequality. Assuming that reverse isoperimetric inequality continues to hold in higher dimensions, and employing the intermediate refined reverse isoperimetric inequalities introduced in [Phys. Rev. Lett. 131, 241401] as guiding principles, we extend the inequality to higher-dimensional spacetimes.
@@
Finally, we consider the Kerr-Sen-AdS black hole. Sen discovered a charged, rotating black hole solution in the low-energy limit of heterotic string theory, now commonly referred to as the Kerr–Sen black hole [40]. The thermodynamics is introduced in the extended phase space in [44, 45]. It can be seen from Eq. (73) in the Supplement that the identity satisfied by the thermodynamic variables is analogous to that of the KN-AdS black hole, with the black hole mass $M$ and thermodynamic volume $V$ replaced by $\widetilde{M}$} and $\widetilde{V}$}, respectively. This is because both the standard form of the first law of black hole thermodynamics and the Bekenstein–Smarr mass formula should be satisfied simultaneously. Consequently, the inequality takes the following form:
$\frac{4\pi J^{2}}{3{\widetilde{M}}{\widetilde{V}}}<1.$ (24)
@@
where $T$, $V$, $\Omega$, and $\Phi$ are the Hawking temperature, thermodynamic volume, angular velocity, and electric potential respectively, we clearly observe that the mass $M$ of the KN-AdS black hole should be essentially identified as enthalpy rather than internal energy. Then, it is straightforward to have
@@
Kerr-Sen-AdS black hole. In the usual Boyer-Lindquist coordinate, the Kerr-Sen-AdS metric reads [44]
$ds^{2}$ $\!\!\!\!\!=\!\!\!\!\!$ $-\frac{\Delta_{r}}{\rho^{2}}\left(dt-\frac{a\sin^{2}\theta}{\Xi}d\phi\right)^{2}+\frac{\rho^{2}}{\Delta_{r}}dr^{2}+\frac{\rho^{2}}{\Delta_{\theta}}d\theta^{2}+\frac{\sin^{2}\theta\Delta_{\theta}}{\rho^{2}}\left(adt-\frac{r^{2}+2br+a^{2}}{\Xi}d\phi\right)^{2},$
where
$\rho^{2}$ $=$ $r^{2}+2br+a^{2}\cos^{2}\theta,\quad\Delta_{r}=(r^{2}+2br+a^{2})\left(1+\frac{r^{2}+2br}{\ell^{2}}\right)-2mr,$
$\Xi$ $=$ $1-\frac{a^{2}}{\ell^{2}},\quad\Delta_{\theta}=1-\frac{a^{2}}{\ell^{2}}\cos^{2}\theta,\quad A=-\frac{qr}{\rho^{2}}\left(dt-\frac{a\sin^{2}\theta}{\Xi}d\phi\right).$ (63)
The parameter $b$ denotes the dilatonic charge of the black holes and is defined by
$b=\frac{q^{2}}{2m},$ (64)
where $q$ and $m$ represent the electric charge and mass of the black holes, respectively. In the limit $\ell\to\infty$, the metric reduces to the conventional Kerr-Sen black holes. For the non-rotating case ($a=0$), it further simplifies to the Gibbons-Maeda-Garfinkle-Horowitz-Strominger (GMGHS) solution. Gibbons and Maeda [41] originally derived the black hole and black brane solutions for the dilaton field, while Garfinkle, Horowitz, and Strominger [42] extended this to include charged configurations. It is worth noting that four-dimensional Kerr-Sen-(A)dS black holes have been studied from multiple perspectives, including their shadows [43] and phase-space thermodynamics in the extended phase space [45].
We are now in a position to study the thermodynamics of the Kerr-Sen-AdS black hole. Within the framework of the extended phase space, the thermodynamic quantities corresponding to the solution (S0.Ex7) can be computed using standard methods and are given by [44]:
$M$ $=$ $\frac{m}{\Xi^{2}},\quad Q=\frac{q}{\Xi},\quad J=\frac{am}{\Xi^{2}},\quad T=\frac{(r_{+}+b)\left(2r_{+}^{2}+4br_{+}+\ell^{2}+a^{2}\right)-m\ell^{2}}{2\pi\left(r_{+}^{2}+2br_{+}+a^{2}\right)\ell^{2}},$
$S$ $=$ $\frac{\pi\left(r_{+}^{2}+2br_{+}+a^{2}\right)}{\Xi},\quad\Omega=\frac{a\Xi}{\left(r_{+}^{2}+2br_{+}+a^{2}\right)},\quad\Phi=\frac{qr_{+}}{\left(r_{+}^{2}+2br_{+}+a^{2}\right)},$ (65)
where the mass term $m$ can be derived from the condition $\Delta_{r}(r=r_{+})=0$ with $r_{+}$ being the radius of the event horizon.
It is straightforward to verify that the thermodynamic quantities given in Eq. (S0.Ex9) satisfy the Bekenstein–Smarr mass relation
$M=2TS+2\Omega J+\Phi Q-2PV,$ (66)
where the thermodynamic volume $V$, defined as [44]
$V=\frac{4\pi}{3\Xi}\left(r_{+}+b\right)\left(r_{+}^{2}+2br_{+}+a^{2}\right)$ (67)
conjugate to the pressure (32). Unfortunately, the first law reduces merely to a differential identity,
$dM=TdS+\Omega dJ+\Phi dQ-VdP+Jd\Xi/(2a).$ (68)
This occurs because the thermodynamic quantities have been defined in a rotating frame at infinity rather than in the rest frame at infinity.
The Kerr-Sen-AdS solution can be transformed into the rest frame at infinity by performing the simple coordinate transformation $\phi=\phi-at/\ell^{2}$. After a straightforward but tedious computation of the thermodynamic quantities in this frame, one finds that only the mass, the angular velocity, and the thermodynamic volume differ from those given in Eq. (S0.Ex9). These quantities are given by [44]
$\widetilde{M}=M+\frac{a}{\ell^{2}}J,\qquad{\widetilde{\Omega}}=\Omega+\frac{a}{\ell^{2}},\qquad{\widetilde{V}}=V+\frac{4\pi}{3}aJ.$ (69)
In this frame, the thermodynamic quantities consistently satisfy both the standard form of the first law of black hole thermodynamics and the Bekenstein–Smarr mass formula simultaneously,
$d\widetilde{M}$ $=$ $TdS+{\widetilde{\Omega}}dJ+\Phi dQ-{\widetilde{V}}dP,$
$\widetilde{M}$ $=$ $2TS+2{\widetilde{\Omega}}J+\Phi Q-2P{\widetilde{V}}$. (70)
It is straightforward to verify that both the differential and integral mass relations above can be derived from the following Christodoulou–Ruffini–type squared-mass formula [44]:
${\widetilde{M}}^{2}=\left(1+\frac{8PS}{3}\right)\left[\left(1+\frac{8PS}{3}\right)\frac{S}{4\pi}+\frac{\pi J^{2}}{S}+\frac{Q^{2}}{2}\right].$ (71)
From the standard first law of thermodynamics in the extended phase space, given in the first line of Eq. (S0.Ex10), we can obtain the thermodynamic volume of the Kerr-Sen-AdS black hole in terms of thermodynamics potentials as
${\widetilde{V}}=\left(\frac{\partial\widetilde{M}}{\partial P}\right)_{S,J,Q}=\frac{4\sqrt{S}\left\{6\pi^{2}J^{2}+S\left[S(8PS+3)+3\pi Q^{2}\right]\right\}}{3\sqrt{\pi}\sqrt{8PS+3}\sqrt{12\pi^{2}J^{2}+S\left[S(8PS+3)+6\pi Q^{2}\right]}}.$ (72)
Solving the equation (72) for pressure and substituting it into the mass formula (71) and using the Bekenstein–Hawking formula $S=A/4$, one can eliminate the pressure and reexpress the mass $\widetilde{M}$} in terms of $(A,{\widetilde{V}},J,Q)$, which leads to the following identity [32]
$36\pi{\widetilde{M}}^{2}{\widetilde{V}}^{2}-{\widetilde{M}}^{2}A^{3}-64\pi^{3}J^{4}=16\pi^{2}Q^{2}J^{2}A.$ (73)
By setting $Q=0$ we find the corresponding identity for uncharged Kerr-Sen-AdS black hole as
$36\pi{\widetilde{M}}^{2}{\widetilde{V}}^{2}-{\widetilde{M}}^{2}A^{3}-64\pi^{3}J^{4}=0.$ (74)